# Echo: Decoupling Inference and Training for Large-Scale RL Alignment on Heterogeneous Swarms

### v2

@@ Abstract
Modern RL-based post-training for large language models (LLMs) co-locate trajectory sampling and policy optimisation on the same GPU cluster, forcing the system to switch between inference and training workloads. This serial context switching violates the single-program–multiple-data (SPMD) assumption underlying today’s distributed training systems. We present Echo, the RL system that cleanly decouples these two phases across heterogeneous “inference” and “training” swarms while preserving statistical efficiency. Echo introduces two lightweight synchronization protocols: a sequential pull mode that refreshes policy weights according to API call for minimal bias, and an asynchronous push–pull mode that streams version-tagged rollouts through a replay buffer to maximise hardware utilisation. Training three representative RL workloads with Qwen3-4B, Qwen2.5-7B, and Qwen3-32B on a geographically distributed cluster, Echo matches a fully co-located Verl baseline in convergence speed and final reward while off-loading trajectory generation to commodity edge hardware. These promising results demonstrate that large-scale RL for LLMs could achieve datacentre-grade performance using decentralised, heterogeneous resources.
@@ 1 Introduction
We train three typical RL workloads from scratch on Qwen3-4B, Qwen2.5-7B, and Qwen3-32B—using both Echo and a fully co-located Verl baseline. In every case, Echo matches the baseline’s convergence speed and final reward while off-loading trajectory generation to a pool of heterogeneous edge devices. These results demonstrate that large-scale RL can exploit decentralized, commodity hardware without sacrificing training efficiency.
@@ 3.1 Experiment Setup
We utilize the open-source Qwen series models as base models for training: Qwen3-4B, Qwen2.5-7B and Qwen3-32B.
@@
3.2 Results
End-to-End Comparison of Training Convergence.
Performance Improvement by Echo.
In the box-pushing task, the Qwen3-4B model post-trained with Echo(GRPO) surpassed the Qwen3-32B, improving by $4\%$. As shown in Table 1. In math reasoning tasks, we post-trained Qwen2.5-7B with Echo on Eurus-2-RL-Math datasetcui2025processreinforcementimplicitrewards , and it outperformed the Qwen2.5-32B on all six test datasets, achieving a $12\%$ improvement on average, as shown in Table 2. Our experiments on the K&K task show that training under Echo significantly improves model performance, especially in more difficult multi-person scenarios (6-8 people). The Qwen3-32B model fine-tuned with GRPO and LoRA achieved near-perfect scores (all metrics $\geq$ 0.99), surpassing o3-mini-high, DeepSeek-R1 and o4-mini, achieving the best performance. See Table 3 for detailed results.
@@ 4 Conclusion and Future Work
We presented Echo, a dual-swarm RL system that disentangles trajectory sampling from policy optimisation so each phase can exploit the hardware that suits it best. Echo combines (i) a fully decentralised, pipeline-parallel inference engine (Parallax) capable of harnessing heterogeneous edge devices and (ii) an enhanced Verl trainer that retains broad algorithm support and production-grade LoRA fine-tuning. Two lightweight synchronisation protocols—sequential for accuracy and asynchronous for efficiency—bound policy lag without requiring datacentre-class interconnects and can be swapped with only minimal orchestration changes. Experiments on Qwen3-4B, Qwen2.5-7B, and Qwen3-32B show that Echo matches a fully co-located baseline in convergence speed and final reward while off-loading trajectory generation to commodity edge hardware, demonstrating the viability of large-scale RL on geographically dispersed resources.

### v3

@@ Abstract
Modern RL-based post-training for large language models (LLMs) co-locate trajectory sampling and policy optimisation on the same GPU cluster, forcing the system to switch between inference and training workloads. This serial context switching violates the single-program–multiple-data (SPMD) assumption underlying today’s distributed training systems. We present Echo, the RL system that cleanly decouples these two phases across heterogeneous “inference” and “training” swarms while preserving statistical efficiency. Echo introduces two lightweight synchronization protocols: a sequential pull mode that refreshes policy weights according to API call for minimal bias, and an asynchronous push–pull mode that streams version-tagged rollouts through a replay buffer to maximise hardware utilisation. Training four representative RL workloads with Qwen3-4B, Qwen2.5-7B, Qwen3-30B-A3B-Thinking-2507, and Qwen3-32B on a geographically distributed cluster, Echo matches a fully co-located Verl baseline in convergence speed and final reward while off-loading trajectory generation to commodity edge hardware. These promising results demonstrate that large-scale RL for LLMs could achieve datacentre-grade performance using decentralised, heterogeneous resources.
@@ 1 Introduction
We train four typical RL workloads from scratch on Qwen3-4B, Qwen2.5-7B, Qwen3-30B-A3B-Thinking-2507, and Qwen3-32B—using both Echo and a fully co-located Verl baseline. In every case, Echo matches the baseline’s convergence speed and final reward while off-loading trajectory generation to a pool of heterogeneous edge devices. These results demonstrate that large-scale RL can exploit decentralized, commodity hardware without sacrificing training efficiency.
@@ 3.1 Experiment Setup
We utilize the open-source Qwen series models as base models for training: Qwen3-4B, Qwen2.5-7B, Qwen3-32B and Qwen3-30B-A3B-Thinking-2507.
@@
Sokoban (Base model: Qwen3-30B-A3B-Thinking-2507): Learning Rate: 1e-6, Batch Size: 8, Rollout N: 16, KL Coefficient: 0.001.
@@
Qwen3-30B-A3B-Thinking-2507 72.75
Qwen3-30B-A3B-Thinking-2507-Echo(GRPO) 82.20
Deepseek-R1 75.75
Qwen3-235B-A22B-Thinking-2507 79.68
gpt-oss-120b 79.69
@@
3.2 Results
End-to-End Comparison of Training Convergence.
Performance Improvement by Echo.
In the Sokoban task, we have post-trained 2 models: Qwen3-4B and Qwen3-30B-A3B-Thinking-2507. As shown in Table 1, Qwen3-4B-Echo surpassed the Qwen3-4B by $13\%$. Noteably, Qwen3-30B-A3B-Thinking-2507-Echo even achieves the best performance over DeepSeek-R1, Qwen3-235B-A22B-Thinking-2507, and gpt-oss-120b. In math reasoning tasks, we post-trained Qwen2.5-7B with Echo on Eurus-2-RL-Math datasetcui2025processreinforcementimplicitrewards , and it outperformed the Qwen2.5-32B on all six test datasets, achieving a $12\%$ improvement on average, as shown in Table 2. Our experiments on the K&K task show that training under Echo significantly improves model performance, especially in more difficult multi-person scenarios (6-8 people). The Qwen3-32B model fine-tuned with GRPO and LoRA achieved near-perfect scores (all metrics $\geq$ 0.99), surpassing o3-mini-high, DeepSeek-R1 and o4-mini, achieving the best performance. See Table 3 for detailed results.
@@ 4 Conclusion and Future Work
We presented Echo, a dual-swarm RL system that disentangles trajectory sampling from policy optimisation so each phase can exploit the hardware that suits it best. Echo combines (i) a fully decentralised, pipeline-parallel inference engine (Parallax) capable of harnessing heterogeneous edge devices and (ii) an enhanced Verl trainer that retains broad algorithm support and production-grade LoRA fine-tuning. Two lightweight synchronisation protocols—sequential for accuracy and asynchronous for efficiency—bound policy lag without requiring datacentre-class interconnects and can be swapped with only minimal orchestration changes. Experiments on Qwen3-4B, Qwen2.5-7B, Qwen3-30B-A3B-Thinking-2507, and Qwen3-32B show that Echo matches a fully co-located baseline in convergence speed and final reward while off-loading trajectory generation to commodity edge hardware, demonstrating the viability of large-scale RL on geographically dispersed resources.